# Data-Centric Foundation Models in Computational Healthcare: A Survey

### v1

@@ 9.4. Regulation
The open-source community is highly valuable for FM development, revealing insights into training data, model architecture, hyperparameter settings, and user interactions. These fine details could help establish strong regulation mechanisms to enhance the trustworthy use of FMs in healthcare. To do so, medical data standardization is essential during the dataset preparation. The standardization will ensure the data collected from multiple institutions are organized in a structured format compatible with FM model development to avoid data inconsistencies. For instance, the Digital Imaging and Communications in Medicine (DICOM) standards and the Picture Archiving and Communication System (PACS) provide a standardized platform for imaging data management. Also, the CEDAR Workbench (Musen et al., 2022) is a web-based platform allowing us to manage and share the library of metadata templates. To enhance the sharing of scientific data, the FAIR (Findable, Accessible, Interoperable, and Reproducible) principle (Wilkinson et al., 2016) provides consistent metadata preparation. Popularizing the standardized guideline documents, such as the ”datasheet” (Gebru et al., 2021), is crucial to simplify metadata preparation. These efforts are helpful for us to understand the comprehensive details of the dataset, enhancing the wide adoption of clinical data to accelerate the progress of FM research in healthcare.
@@ Appendix A Healthcare Data Modalities
Histopathology involves examining visually perceptible changes in cells and tissues to study the manifestations of diseases (Kumar et al., 2020, 2019). In clinical practice, tissue is first removed from the body, and then the histological sections are placed onto glass slides. A high-resolution slide scanner equipped with microscope optics serially captures image tiles across the tissue section. The seamless reconstruction of these tile images yields a single monolithic digital representation of the entire glass slide, typically within the order of billions of pixels. Staining techniques highlight architectural and morphological details, demonstrating the changes in cells and the actual causes of the illness. Therefore, the microscopic examination of stained tissue sections offers a definitive disease diagnosis. Different magnifications in histopathology imaging can reveal morphological details, offering multi-level visual perceptions in clinical scenarios. Higher magnifications reveal finer attributes, while lower magnifications facilitate the understanding of overall tissue morphology (Ashtaiwi, 2022). Due to the high resolution of histopathological images, current FMs mostly take split image patches as input instead of whole slides (Huang et al., 2023a; Lu et al., 2023a).
@@ Appendix B Healthcare Foundation Models
Table 2. Typical foundation models in healthcare and medicine.
@@
Model Base Initial Model Weights Pre-training Data SciBERT (Beltagy et al., 2019) BERT (Devlin et al., 2018) - Semantic Scholar (Ammar et al., 2018) Clinical BERT (Alsentzer et al., 2019) BERT BERT MIMIC-III (Johnson et al., 2016) BlueBERT (Peng et al., 2019) BERT - PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/ + MIMIC-III BioBERT (Lee et al., 2020) BERT BERT PubMed + PMC [^7] https://www.ncbi.nlm.nih.gov/pmc/ PubMedBERT (Gu et al., 2021) BERT - PubMed BioLinkBERT (Yasunaga et al., 2022) BERT - PubMed BioGPT (Luo et al., 2022) GPT-2 (Radford et al., 2019) - PubMed Clinical-T5 (Lehman and Johnson, 2023) T5 (Raffel et al., 2020) T5 MIMIC-III + MIMIC-IV (Johnson et al., 2020) HIPT (Chen et al., 2022) DINO (Caron et al., 2021) - TCGA [^8] https://www.cancer.gov/ccg/research/genome-sequencing/tcga CTransPath (Wang et al., 2022d) SRCL (Wang et al., 2022d) - TCGA + PAIP [^9] http://www.wisepaip.org/paip/ RETFound (Zhou et al., 2023a) MAE (He et al., 2022) Vision Transformer (Dosovitskiy et al., 2020) RETFound REMEDIS (Azizi et al., 2023) SimCLR (Chen et al., 2020) BiT (Beyer et al., 2022) MIMIC-IV + CheXpert (Irvin et al., 2019) Virchow (Vorontsov et al., 2023) DINOv2 (Oquab et al., 2023) - Virchow PubMedCLIP (Eslami et al., 2021) CLIP (Radford et al., 2021) CLIP ROCO (Pelka et al., 2018) CheXzero (Tiu et al., 2022) CLIP CLIP MIMIC-CXR (Johnson et al., 2019) MedCLIP (Wang et al., 2022c) CLIP Clinical BERT + SwinTransformer (Liu et al., 2021) CheXpert + MIMIC-CXR BiomedCLIP (Zhang et al., 2023e) CLIP PubMedBERT PMC-15M (Zhang et al., 2023e) PMC-CLIP (Lin et al., 2023b) CLIP PubMedBERT PMC-OA (Lin et al., 2023b) MI-Zero (Lu et al., 2023b) CLIP HistPathGPT (Lu et al., 2023b) + CTransPath (Wang et al., 2022e) ARCH (Gamper and Rajpoot, 2021) PLIP (Huang et al., 2023a) CLIP CLIP OpenPath (Huang et al., 2023a) QuiltNet (Ikezogwo et al., 2023) CLIP CLIP Quilt-1M (Ikezogwo et al., 2023) KAD (Zhang et al., 2023c) CLIP - MIMIC-CXR + UMLS (Bodenreider, 2004) CONCH (Lu et al., 2023a) CoCa (Yu et al., 2022) - PubMed + PMC Med-Flamingo (Moor et al., 2023b) Flamingo (Alayrac et al., 2022) Flamingo MTB (Moor et al., 2023b) + PMC-OA Med-PaLM (Singhal et al., 2022) PaLM (Chowdhery et al., 2022) PaLM MultiMedQA (Singhal et al., 2022) ChatDoctor (Yunxiang et al., 2023) LLaMA (Touvron et al., 2023a) LLaMA HealthCareMagic-100k (Yunxiang et al., 2023) PMC-LLaMA (Wu et al., 2023d) LLaMA LLaMA MedC (Wu et al., 2023d) Clinical LLaMA-LoRA (Gema et al., 2023) LLaMA LLaMA / PMC-LLaMA MIMIC-IV (Johnson et al., 2020) BioMedGPT (Luo et al., 2023) LLaMA 2 (Touvron et al., 2023b) LLaMA 2 S2ORC (Lo et al., 2019) LLaVA-Med (Li et al., 2023e) LLaVA (Liu et al., 2023b) LLaVA PMC-15M (Zhang et al., 2023e) + GPT-4 (OpenAI, 2023) Qilin-Med-VL (Liu et al., 2023c) LLaVA Chinese-LLaMA2 + CLIP ChiMed-VL (Liu et al., 2023c) MedVInT (Zhang et al., 2023d) - PMC-LLaMA + PMC-CLIP PMC-VQA (Zhang et al., 2023d) RadFM (Wu et al., 2023e) - PMC-LLaMA MedMD (Wu et al., 2023e) ESM-2 (Lin et al., 2023a) Transformer (Vaswani et al., 2017) - UniRef (Suzek et al., 2015) GET (Fu et al., 2023) Transformer - GET AlphaMissense (Cheng et al., 2023a) AlphaFold (Jumper et al., 2021) - PDB (Berman et al., 2000) + UniRef

### v2

@@ 9.4. Regulation
The open-source community is highly valuable for FM development, revealing insights into training data, model architecture, hyperparameter settings, and user interactions. These fine details could help establish strong regulation mechanisms to enhance the trustworthy use of FMs in healthcare. To do so, medical data standardization is essential during the dataset preparation. The standardization will ensure the data collected from multiple institutions are organized in a structured format compatible with FM model development to avoid data inconsistencies. For instance, the Digital Imaging and Communications in Medicine (DICOM) standards and the Picture Archiving and Communication System (PACS) provide a standardized platform for imaging data management. Also, the CEDAR Workbench (Musen et al., 2022) is a web-based platform allowing us to manage and share the library of metadata templates. To enhance the sharing of scientific data, the FAIR (Findable, Accessible, Interoperable, and Reproducible) principle (Wilkinson et al., 2016) provides consistent metadata preparation. Popularizing the standardized guideline documents, such as the ”datasheet” (Gebru et al., 2021), is crucial to simplify metadata preparation. These efforts are helpful for us to understand the comprehensive details of the dataset, which in turn promotes the wide adoption of clinical data to advance FM research in healthcare.
@@ Appendix A Healthcare Data Modalities
Histopathology involves examining visually perceptible changes in cells and tissues to study the manifestations of diseases (Kumar et al., 2020, 2019). In clinical practice, tissue is first removed from the body, and then the histological sections are placed onto glass slides. A high-resolution slide scanner equipped with microscope optics serially captures image tiles across the tissue section. The seamless reconstruction of these tile images yields a single monolithic digital representation of the entire glass slide, typically within the order of billions of pixels. Staining techniques highlight architectural and morphological details, demonstrating the changes in cells and the actual causes of the illness. Therefore, the microscopic examination of stained tissue sections offers a definitive disease diagnosis. Different magnifications in histopathology imaging can reveal morphological details, offering multi-level visual perceptions in clinical scenarios. Higher magnifications reveal finer attributes, while lower magnifications facilitate the understanding of overall tissue morphology (Ashtaiwi, 2022). Due to the high resolution of histopathological images, current FMs mostly take split image patches as input instead of whole slides (Huang et al., 2023a; Lu et al., 2023b).
@@ Appendix B Healthcare Foundation Models
Table 2. Typical foundation models in healthcare and medicine. A star (*) after the pre-training data shows that the authors constructed the data with more than two sources.
@@
Model Base Initial Model Weights Pre-training Data SciBERT (Beltagy et al., 2019) BERT (Devlin et al., 2018) - Semantic Scholar (Ammar et al., 2018) Clinical BERT (Alsentzer et al., 2019) BERT BERT MIMIC-III (Johnson et al., 2016) BlueBERT (Peng et al., 2019) BERT - PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/ + MIMIC-III BioBERT (Lee et al., 2020) BERT BERT PubMed + PMC [^7] https://www.ncbi.nlm.nih.gov/pmc/ PubMedBERT (Gu et al., 2021) BERT - PubMed BioLinkBERT (Yasunaga et al., 2022) BERT - PubMed BioGPT (Luo et al., 2022) GPT-2 (Radford et al., 2019) - PubMed Med-PaLM (Singhal et al., 2022) PaLM (Chowdhery et al., 2022) PaLM MedQA (Jin et al., 2021) Clinical-T5 (Lehman and Johnson, 2023) T5 (Raffel et al., 2020) T5 MIMIC-III + MIMIC-IV (Johnson et al., 2020) ChatDoctor (Yunxiang et al., 2023) LLaMA (Touvron et al., 2023a) LLaMA HealthCareMagic [^8] https://www.askadoctor24x7.com/ PMC-LLaMA (Wu et al., 2023d) LLaMA LLaMA MedC (Wu et al., 2023d) Med-PaLM 2 (Singhal et al., 2023b) PaLM 2 - MedQA Clinical LLaMA-LoRA (Gema et al., 2023) LLaMA LLaMA / PMC-LLaMA MIMIC-IV BioMedGPT (Luo et al., 2023) LLaMA 2 (Touvron et al., 2023b) LLaMA 2 S2ORC (Lo et al., 2019) HIPT (Chen et al., 2022) DINO (Caron et al., 2021) - TCGA [^9] https://www.cancer.gov/ccg/research/genome-sequencing/tcga CTransPath (Wang et al., 2022d) SRCL (Wang et al., 2022d) - TCGA + PAIP [^10] http://www.wisepaip.org/paip/ RETFound (Zhou et al., 2023a) MAE (He et al., 2022) Vision Transformer (Dosovitskiy et al., 2020) * Virchow (Vorontsov et al., 2023) DINOv2 (Oquab et al., 2023) - * REMEDIS (Azizi et al., 2023) SimCLR (Chen et al., 2020) BiT (Beyer et al., 2022) MIMIC-IV + CheXpert (Irvin et al., 2019) UNI (Chen et al., 2023b) DINOv2 - Mass-100K* (Chen et al., 2023b) RudolfV (Dippel et al., 2024) DINOv2 DINOv2 * Pai et al. (Pai et al., 2024) SimCLR - * Prov-GigaPath (Xu et al., 2024) MAE - Prov-Path* (Xu et al., 2024) PubMedCLIP (Eslami et al., 2021) CLIP (Radford et al., 2021) CLIP ROCO (Pelka et al., 2018) CheXzero (Tiu et al., 2022) CLIP CLIP MIMIC-CXR (Johnson et al., 2019) MedCLIP (Wang et al., 2022c) CLIP Clinical BERT + SwinTransformer (Liu et al., 2021) CheXpert + MIMIC-CXR BiomedCLIP (Zhang et al., 2023e) CLIP PubMedBERT PMC-15M* (Zhang et al., 2023e) PMC-CLIP (Lin et al., 2023b) CLIP PubMedBERT PMC-OA* (Lin et al., 2023b) MedVInT (Zhang et al., 2023d) - PMC-LLaMA + PMC-CLIP PMC-VQA* (Zhang et al., 2023d) LLaVA-Med (Li et al., 2023e) LLaVA (Liu et al., 2023b) LLaVA PMC-15M (Zhang et al., 2023e) + GPT-4 (OpenAI, 2023) MI-Zero (Lu et al., 2023c) CLIP HistPathGPT (Lu et al., 2023c) + CTransPath (Wang et al., 2022e) ARCH (Gamper and Rajpoot, 2021) PLIP (Huang et al., 2023a) CLIP CLIP OpenPath* (Huang et al., 2023a) QuiltNet (Ikezogwo et al., 2023) CLIP CLIP Quilt-1M* (Ikezogwo et al., 2023) CONCH (Lu et al., 2023b) CoCa (Yu et al., 2022) - PubMed + PMC Med-Flamingo (Moor et al., 2023b) Flamingo (Alayrac et al., 2022) Flamingo MTB (Moor et al., 2023b) + PMC-OA KAD (Zhang et al., 2023c) CLIP - MIMIC-CXR + UMLS (Bodenreider, 2004) RadFM (Wu et al., 2023e) - PMC-LLaMA MedMD* (Wu et al., 2023e) Qilin-Med-VL (Liu et al., 2023d) LLaVA Chinese-LLaMA2 + CLIP ChiMed-VL* (Liu et al., 2023d) PathChat (Lu et al., 2023a) LLaVA LLaMA 2 + UNI PathChatInstruct* (Lu et al., 2023a) CheXagent (Chen et al., 2024) BLIP-2 (Li et al., 2023c) Mistral 7B (Jiang et al., 2023) CheXinstruct* (Chen et al., 2024) EchoCLIP (Christensen et al., 2024) CLIP CLIP * PRISM (Shaikovski et al., 2024) CoCa BioGPT + Virchow * Med-Gemini (Saab et al., 2024) Gemini (Reid et al., 2024) Gemini * RadFound (Liu et al., 2024b) - - RadVLCorpus* (Liu et al., 2024b) ESM-2 (Lin et al., 2023a) Transformer (Vaswani et al., 2017) - UniRef (Suzek et al., 2015) GET (Fu et al., 2023) Transformer - * AlphaMissense (Cheng et al., 2023a) AlphaFold (Jumper et al., 2021) - PDB (Berman et al., 2000) + UniRef MoleculeSTM (Liu et al., 2023c) CLIP - PubChem (Kim et al., 2023)